# A bound on the quantum value of all compiled nonlocal games

### v1

@@ 5.3. Quantum strategies and correlations
A strongly non-signaling algebraic strategy for $\mathcal{G}_{\textnormal{seq}}$ consists of
@@
Any quantum strategy for $\mathcal{G}$ can be converted into such an algebraic strategy, provided it satisfies the strong non-signaling property for a $C^{*}$-algebra $\mathscr{B}$. Simply define the positive linear functionals $\phi_{xa}$ by $\phi_{xa}(B):=\operatorname{tr}\left(\sigma_{xa}B\right)$ for all $B\in\mathscr{B}$. Thus the algebraic model is at least as general.
@@
For any strongly non-signaling algebraic strategy, the resulting correlation is a nonlocal commuting operator correlation, that is, in $C_{qc}$.
In the proof of the theorem we will use the following version of the Radon-Nikodyme theorem for $C^{*}$-algebras, which is well-known to experts in operator algebras. We refer the reader to Section 2.2 for the concepts used in its statement.
Proposition 5.15 (Radon-Nikodyme theorem for $C^{*}$-algebras).
@@
The version stated here is [Bla06, Prop. II.6.4.6] and we refer to this reference for a concise proof.
@@
Observe that $\phi_{xa}\leq\phi$ for all $x\in\mathcal{I}_{A}$ and $a\in\mathcal{O}_{A}$. Let $(\mathcal{H}_{\phi},\pi_{\phi},\lvert\nu_{\phi}\rangle)$ be a GNS triple associated with $\phi$. Then, by Proposition 5.15, for each pair $(x,a)$ there exists an operator $M_{xa}\in\pi_{\phi}(\mathscr{B})^{\prime}$ (in the commutant) such that $0\leq M_{xa}\leq\mathbbm{1}$ and we have $\phi_{xa}(B)=\langle\nu_{\phi}\rvert M_{xa}\pi_{\phi}(B)\lvert\nu_{\phi}\rangle$ for all $B\in\mathscr{B}$. Because $B_{yb}\in\mathscr{B}$ for all $y,b$, it follows that, for all $x,y,a,b$,
@@
$=\phi(E^{*}F)=\langle\nu_{\phi}\rvert\phi(E^{*})\phi(F)\lvert\nu_{\phi}\rangle$
where we used that $M_{xa}\in\pi_{\phi}(\mathscr{B})^{\prime}$ and that $\pi_{\phi}$ is a *-homomorphism. Thus we have $\langle\nu_{\phi}\rvert\pi_{\phi}(E^{*})(\sum_{a\in\mathcal{O}_{A}}M_{xa}-\mathbbm{1})\pi_{\phi}(F)\lvert\nu_{\phi}\rangle=0$ for all $E,F\in\mathscr{B}$. Since $\pi_{\phi}(\mathscr{B})\lvert\nu_{\phi}\rangle$ is dense in $\mathcal{H}_{\phi}$, we deduce that $\sum_{a\in\mathcal{O}_{A}}M_{xa}=\mathbbm{1}$, as desired, concluding the proof. ∎
Finally, any commuting operator strategy for the nonlocal game (Definition 3.5) gives rise to a strongly non-signaling quantum strategy for the sequential game. Simply use the same Hilbert space $\mathcal{H}$, $\sigma_{xa}=\sqrt{M_{xa}}\lvert\psi\rangle\langle\mathopen{}\psi\rvert\sqrt{M_{xa}}$, the operators $B_{yb}=N_{yb}$, and let $\mathscr{B}$ denote the $C^{*}$-algebra generated by the operators $\{B_{yb}\}_{b\in\mathcal{O}_{B},y\in\mathcal{I}_{B}}$. Then it is easily verified that strong non-signaling holds for $\mathscr{B}$, noting that $M_{xa}\in\mathscr{B}^{\prime}$ (as we started from a commuting operator strategy) and hence the same is true for its positive square roots $\sqrt{M_{xa}}$. Altogether we obtain the following corollary which may be of independent interest.
@@
The commuting operator correlation set $C_{qc}$ is equal to the correlations produced by strongly non-signaling algebraic sequential strategies (Definition 5.13), as well as to the correlations produced by (possibly infinite-dimensional) strongly non-signaling quantum sequential strategies (Definition 5.9).
Remark 5.17.
Clearly, a quantum strategy is strongly non-signaling (Definition 5.9) if $\sigma_{x}=\sigma_{x^{\prime}}$ for all $x,x^{\prime}\in\mathcal{I}_{A}$. This condition is analogous to the one that we gave in the classical case. If $\mathcal{H}$ is finite-dimensional then we can always reduce to this situation, by replacing each $\sigma_{xa}$ by its twirl over the unitary group of the commutant of the $C^{*}$-algebra generated by the POVM elements. When $\mathcal{H}$ is infinite-dimensional such a reduction is not possible, but the $C^{*}$-algebraic model described in Definition 5.13 can serve as a useful substitute. We note that when $\mathscr{B}$ is finite-dimensional and commutative, the latter reduces to the situation of Proposition 5.4.
6. Upper bound on the quantum value of compiled nonlocal games
In this section we prove that the (maximal) quantum value of a compiled game never exceeds the commuting operator value of the corresponding nonlocal game. The basic idea is as follows. In Section 4, we showed that any QPT strategy of a nonlocal game satisfies an analogue of the strong non-signaling condition discussed in Section 5. More precisely, Proposition 4.6 states that (5.2) holds to arbitrary precision when the security parameter tends to infinity, for any fixed polynomial in the Bob POVMs. We would like to take a limit, but as the Hilbert spaces will depend on the security parameter, we instead work with a single universal $C^{*}$-algebra. We can then define a sequence of states on this algebra, which captures precisely all information that can be accessed using the Bob POVMs, for every value of the security parameter, and use compactness of the state space of a $C^{*}$-algebra to define a limit where the strong non-signaling condition holds exactly. The result then follows from Theorem 5.14.
We now describe the required $C^{*}$-algebra, which is often called the POVM algebra, denoted $\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I},\mathcal{O}}$, for finite sets $\mathcal{I}$ and $\mathcal{O}$ [PSZZ23]. It has elements $\{B_{xa}\}_{x\in\mathcal{I},a\in\mathcal{O}}$ which satisfy the relations $0\leq B_{xa}\leq\mathbbm{1}$ and $\sum_{a\in\mathcal{O}}B_{xa}=1$ for each $x\in\mathcal{I}$. Importantly, it satisfies the following universal property: for any Hilbert space $\tilde{\mathcal{H}}$ and any collection of POVMs $\{\tilde{B}_{xa}\}$ on $\tilde{\mathcal{H}}$, there exists a unique *-homomorphism $\theta\colon\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I},\mathcal{O}}\to\mathcal{B}(\tilde{\mathcal{H}})$ sending $B_{xa}\mapsto\tilde{B}_{xa}$ for all $x\in\mathcal{I}$ and $a\in\mathcal{O}$. The POVM $C^{*}$-algebras are separable as they are finitely generated.
@@ 6. Upper bound on the quantum value of compiled nonlocal games
(6.1) $\phi_{xa}^{\lambda}\colon\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}\to\mathbb{C},\quad\phi_{xa}^{\lambda}(\cdot)=\operatorname{tr}(\sigma_{xa}^{\lambda}\vartheta_{\lambda}(\cdot)).$
@@ Acknowledgements
CP thanks the hospitality at RUB for a productive visit. CP acknowledges funding support from the Natural Sciences and Engineering Research Council of Canada (NSERC). AK, GM, SS and MW acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972. GM is also supported by the European Research Council through an ERC Starting Grant (Grant agreement No. 101077455, ObfusQation). MW is supported by the European Research Council through an ERC Starting Grant (Grant agreement No. 101040907, SYMOPTIC), by the NWO through grant OCENW.KLEIN.267, and by the BMBF through project Quantum Methods and Benchmarks for Resource Allocation (QuBRA).

### v3

@@ 5.4 A characterization of quantum commuting correlations
A strongly non-signaling algebraic strategy consists of
@@
Any quantum strategy for $\mathcal{G}$ can be converted into such an algebraic strategy provided it satisfies the strong non-signaling property with respect to any $C^{*}$-algebra $\mathscr{B}$. Simply define the positive linear functionals $\phi_{xa}$ by $\phi_{xa}(B):=\operatorname{tr}\left(\sigma_{xa}B\right)$ for all $B\in\mathscr{B}$. Thus the algebraic model is at least as general.
@@
For any strongly non-signaling algebraic strategy, the resulting correlation is a (nonlocal) commuting operator correlation, that is, in $C_{qc}$.
In the proof of the theorem we will use the following version of the Radon-Nikodym theorem for $C^{*}$-algebras, which is well-known to experts in operator algebras. We refer the reader to Section 2.2 for the concepts used in its statement.
Proposition 5.16 (Radon-Nikodym theorem for $C^{*}$-algebras).
@@
The version stated here is [5, Prop. II.6.4.6] and we refer to this reference for a concise proof.
@@
Observe that $\phi_{xa}\leq\phi$ for all $x\in\mathcal{I}_{A}$ and $a\in\mathcal{O}_{A}$. Let $(\mathcal{H}_{\phi},\pi_{\phi},\lvert\nu_{\phi}\rangle)$ be a GNS triple associated with $\phi$. Then, by Proposition 5.16, for each pair $(x,a)$ there exists an operator $M_{xa}\in\pi_{\phi}(\mathscr{B})^{\prime}$ (in the commutant) such that $0\leq M_{xa}\leq\mathbbm{1}$ and we have, for all $B\in\mathscr{B}$,
$\phi_{xa}(B)=\langle\nu_{\phi}\rvert M_{xa}\pi_{\phi}(B)\lvert\nu_{\phi}\rangle.$ (5.3)
Because $B_{yb}\in\mathscr{B}$ for all $y,b$, it follows that, for all $x,y,a,b$,
@@
$=\phi(E^{*}F)=\langle\nu_{\phi}\rvert\pi_{\phi}(E^{*})\pi_{\phi}(F)\lvert\nu_{\phi}\rangle$
where we used that $M_{xa}\in\pi_{\phi}(\mathscr{B})^{\prime}$ and that $\pi_{\phi}$ is a *-homomorphism. Thus we have
$\langle\nu_{\phi}\rvert\pi_{\phi}(E^{*})(\sum_{a\in\mathcal{O}_{A}}M_{xa}-\mathbbm{1})\pi_{\phi}(F)\lvert\nu_{\phi}\rangle=0$
for all $E,F\in\mathscr{B}$. Since $\pi_{\phi}(\mathscr{B})\lvert\nu_{\phi}\rangle$ is dense in $\mathcal{H}_{\phi}$, we deduce that $\sum_{a\in\mathcal{O}_{A}}M_{xa}=\mathbbm{1}$, as desired, concluding the proof. ∎
Finally, any commuting operator strategy for the nonlocal game (Definition 3.5) gives rise to a strongly non-signaling quantum strategy. Simply use the same Hilbert space $\mathcal{H}$, $\sigma_{xa}=\sqrt{M_{xa}}\lvert\psi\rangle\langle\mathopen{}\psi\rvert\sqrt{M_{xa}}$, the operators $B_{yb}=N_{yb}$, and let $\mathscr{B}$ denote the $C^{*}$-algebra generated by the operators $\{B_{yb}\}_{b\in\mathcal{O}_{B},y\in\mathcal{I}_{B}}$. Then it is easily verified that strong non-signaling holds for $\mathscr{B}$, noting that $M_{xa}\in\mathscr{B}^{\prime}$ (as we started from a commuting operator strategy) and hence the same is true for its positive square roots $\sqrt{M_{xa}}$. Altogether we obtain the following corollary which may be of independent interest.
@@
The commuting operator correlation set $C_{qc}$ is equal to the correlations produced by strongly non-signaling algebraic strategies (Definition 5.14), as well as to the correlations produced by (possibly infinite-dimensional) strongly non-signaling quantum sequential strategies (Definition 5.9).
As mentioned in Section 1.2, this characterization of the quantum commuting operator correlations was independently established in the context of steering [1, Cor. 5.3].
6 Upper bound on the quantum value of compiled nonlocal games
In this section we prove that the quantum value of a compiled game never exceeds the commuting operator value of the corresponding nonlocal game (Section 6.1). Using the same techniques, we also deduce a self-testing result (Section 6.2)
6.1 Upper bound on the quantum value
The basic idea is as follows. In Section 4, we showed that any QPT strategy of a nonlocal game satisfies an analogue of the strong non-signaling condition discussed in Section 5. More precisely, Proposition 4.6 states that (5.2) holds to arbitrary precision when the security parameter tends to infinity, for any fixed polynomial in the Bob POVMs. We would like to take a limit, but as the Hilbert spaces will depend on the security parameter, we instead work with a single universal $C^{*}$-algebra. We can then define a sequence of states on this algebra, which captures precisely all information that can be accessed using the Bob POVMs, for every value of the security parameter, and use compactness of the state space of a $C^{*}$-algebra to define a limit where the strong non-signaling condition holds exactly. The result then follows from Theorem 5.15. As an application of Theorem 6.1 we state a result concerning commuting operator self-testing for compiled games.
We now describe the required $C^{*}$-algebra, denoted $\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$ for finite sets $\mathcal{I}_{B}$ and $\mathcal{O}_{B}$, which is often called the POVM algebra [38]. It has elements $\{B_{yb}\}_{y\in\mathcal{I}_{B},b\in\mathcal{O}_{B}}$ which satisfy the relations $0\leq B_{yb}\leq\mathbbm{1}$ and $\sum_{b\in\mathcal{O}_{B}}B_{yb}=1$ for each $y\in\mathcal{I}_{B}$. Importantly, it satisfies the following universal property: for any Hilbert space $\tilde{\mathcal{H}}$ and any collection of POVMs $\{\tilde{B}_{yb}\}$ on $\tilde{\mathcal{H}}$, there eyists a unique *-homomorphism $\theta\colon\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}\to\mathcal{B}(\tilde{\mathcal{H}})$ sending $B_{yb}\mapsto\tilde{B}_{yb}$ for all $y\in\mathcal{I}_{B}$ and $b\in\mathcal{O}_{B}$. The POVM $C^{*}$-algebras are separable as they are finitely generated.
@@ 6.1 Upper bound on the quantum value
$\phi_{xa}^{\lambda}\colon\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}\to\mathbb{C},\quad\phi_{xa}^{\lambda}(\cdot)=\operatorname{tr}(\sigma_{xa}^{\lambda}\vartheta_{\lambda}(\cdot)).$ (6.1)
@@
6.2 Self-testing in compiled nonlocal games
Many applications of nonlocal games take advantage of the concept of self-testing; prominent examples include [40, 12, 21]. The standard definition is as follows: a self-test is a quantum correlation such that any quantum strategy realizing the correlation is equivalent to an ideal strategy via local dilations (isometries). In particular, this implies that in any quantum strategy realizing those correlations, the expectation value of any polynomial in Alice’s and Bob’s POVM operators is uniquely determined. This point of view suggests a more abstract definition of self-testing, which has recently been described in [38]. It has the advantage that it can also be adapted to the commuting operator setting.
To describe their definition, we first observe that any commuting operator strategy (Definition 3.5) gives rise to a state $\Psi$ on the $C^{*}$-algebra $\mathscr{A}:=\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{A},\mathcal{O}_{A}}\otimes_{\max}\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$, via the universal property. Here, $\otimes_{\max}$ denotes the max tensor product of the two POVM algebras, whose generators we denote by $\{A_{xa}\}$ and $\{B_{yb}\}$, respectively (see Section 6.1 for their definition). Concretetly, this state is uniquely defined by the property that, for any noncommutative polynomial $P$,
$\Psi(P(\{A_{xa},B_{yb}\}))=\langle\psi\rvert P(\{M_{xa},N_{yb}\})\lvert\psi\rangle.$ (6.3)
The converse also holds: for any state on $\mathscr{A}$ we can obtain a commuting operator strategy, by the GNS construction. Hence, a correlation $p(a,b|x,y)$ is a commuting operator correlation if, and only if, there exists a state $\Psi$ on $\mathscr{A}$ such that $\Psi(A_{xa}\otimes B_{yb})=p(a,b|x,y)$ for all $a\in\mathcal{O}_{A},b\in\mathcal{O}_{B},x\in\mathcal{I}_{A}$, $y\in\mathcal{I}_{B}$. This characterization is well known [23, 18, 37]. The insight of [38] is that this gives rise to a natural definition of self-testing. For example, a commuting operator self-test is simply a correlation $p$ such that there is a unique such state $\Psi$ [38, Def. 7.1]. [^8] Similarly, the finite-dimensional states on $\mathscr{A}$ (i.e., those for which the GNS Hilbert space is finite-dimensional) correspond precisely to the quantum correlations, and one recovers the traditional definition of self-testing by demanding that there is a unique finite-dimensional state realizing the given correlations. Very often, self-tests arise from nonlocal games. We make the following definition:
Definition 6.3.
A nonlocal game $\mathcal{G}$ is called a commuting operator self-test if any commuting operator strategy $S$ such that $\omega(\mathcal{G},S)=\omega_{qc}(\mathcal{G})$ determines the same state $\Psi$ on $\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{A},\mathcal{O}_{A}}\otimes_{\max}\mathscr{A}_{\textnormal{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$ via Eq. 6.3. We say that $\mathcal{G}$ is a commuting operator self-test with corresponding state $\Psi$ to indicate $\Psi$ explicitly.
While less well-known than ordinary self-tests, there are many examples of commuting operator self-tests. For instance, the CHSH game is not just an ordinary self-test but even a commuting operator self-test, meaning that any optimal commuting operator strategy gives rise to the same (finite-dimensional) state [38, 17].
To derive our self-testing result, it will be useful to translate this definition into the language of sequential games:
Lemma 6.4.
Let $\mathcal{G}$ be a commuting operator self-test with corresponding state $\Psi$. Let $\tilde{S}$ be a strongly non-signaling algebraic strategy for $\mathcal{G}_{\textnormal{seq}}$, with positive linear functionals $\phi_{xa}$ and POVMs $\{\tilde{B}_{yb}\}$. If $\omega(\mathcal{G}_{\textnormal{seq}},\tilde{S})=\omega_{qc}(\mathcal{G})$, then it holds, for all $x\in\mathcal{I}_{A}$ and $a\in\mathcal{O}_{A}$ and for every noncommutative polynomial $P$, that
$\phi_{xa}(P(\{\tilde{B}_{yb}\}))=\Psi(A_{xa}\otimes P(\{B_{yb}\})).$
Proof.
Let $\mathscr{B}$ denote the $C^{*}$-algebra generated by the POVM elements $\{\tilde{B}_{yb}\}$, so that $\tilde{S}$ is strongly non-signaling with respect to $\mathscr{B}$. Let $(\mathcal{H}_{\phi},\pi_{\phi},\lvert\nu_{\phi}\rangle)$ be a GNS triple associated with $\phi=\sum_{a}\phi_{xa}$ (which is independent of $x$). In Eq. 5.3 in the proof of Theorem 5.15 we showed that for any strongly non-signaling algebraic strategy there are POVMs $\{M_{xa}\}$ in $\pi_{\phi}(\mathscr{B})^{\prime}$ such that, for all $B\in\mathscr{B}$, $\phi_{xa}(B)=\langle\nu_{\phi}\rvert M_{xa}\pi_{\phi}(B)\lvert\nu_{\phi}\rangle$. In particular, setting $N_{yb}:=\pi_{\phi}(B_{yb})$, we have we have
$\phi_{xa}(P(\{\tilde{B}_{yb}\}))=\langle\nu_{\phi}\rvert M_{xa}\pi_{\phi}(P(\{B_{yb}\}))\lvert\nu_{\phi}\rangle=\langle\nu_{\phi}\rvert M_{xa}P(\{N_{yb}\})\lvert\nu_{\phi}\rangle$
and $[M_{xa},N_{by}]=0$ for all $x\in\mathcal{I}_{A},y\in\mathcal{I}_{B},a\in\mathcal{O}_{A},b\in\mathcal{O}_{B}$. Thus we have constructed a commuting operator $S$ strategy that in particular produces the same correlations, $\phi_{xa}(\tilde{B}_{yb})=\langle\nu_{\phi}\rvert M_{xa}N_{yb}\lvert\nu_{\phi}\rangle$, and hence $\omega(\mathcal{G},S)=\omega(\mathcal{G}_{\textnormal{seq}},\tilde{S})=\omega_{qc}(\mathcal{G})$. Because $\mathcal{G}$ is a commuting operator self-test, we must have
$\langle\nu_{\phi}\rvert M_{xa}P(\{N_{yb}\})\lvert\nu_{\phi}\rangle=\Psi(A_{xa}\otimes P(\{B_{yb}\})),$
which concludes the proof. ∎
We now provide the asymptotic self-testing statement.
Theorem 6.5.
Let $\mathcal{G}$ be a commuting operator self-test with corresponding state $\Psi$. If $S$ is a QPT strategy for the compiled game such that $\lim_{\lambda\to\infty}\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S)=\omega_{qc}(\mathcal{G})$, then it holds that
$\lim_{\lambda\to\infty}\operatorname{tr}\left(\sigma_{xa}^{\lambda}\,P(\{B^{\lambda}_{yb}\})\right)=\Psi(A_{xa}\otimes P(\{B_{yb}\}))$
for every $x\in\mathcal{I}_{A},a\in\mathcal{O}_{A}$ and for every noncommutative polynomial $P$. In particular:
$\lim_{\lambda\to\infty}\operatorname{tr}\left(\sigma_{x}^{\lambda}\,P(\{B^{\lambda}_{yb}\})\right)=\Psi(P(\{B_{yb}\})).$
Proof.
To show the desired convergence, it suffices to show that for any fixed $x$, $a$, $P$ and for any subsequence $\{\lambda_{\ell}\}_{\ell\in\mathbb{N}}$ there is a further subsequence $\{\lambda_{\ell_{k}}\}_{k\in\mathbb{N}}$ such that
$\lim_{k\to\infty}\operatorname{tr}\left(\sigma_{xa}^{\lambda_{\ell_{k}}}\,P(\{B^{\lambda_{\ell_{k}}}_{yb}\})\right)=\Psi(A_{xa}\otimes P(\{B_{yb}\}))$ (6.4)
We will show that for any subsequence $\{\lambda_{\ell}\}_{\ell\in\mathbb{N}}$ there in fact exists a further subsequence $\{\lambda_{\ell_{k}}\}_{k\in\mathbb{N}}$ such that the above holds for all $x,a,P$. To this end, note that our assumption implies that $\lim_{\ell\to\infty}\omega_{\lambda_{\ell}}(\mathcal{G}_{\textnormal{comp}},S)=\omega_{qc}(\mathcal{G})$. Proceeding like in the proof of Theorem 6.1, but with the subsequence $\{\lambda_{\ell}\}$ in place of $\{\lambda\}$, we obtain a further subsequence $\{\lambda_{\ell_{k}}\}_{k\in\mathbb{N}}$ and a strongly non-signaling algebraic strategy $\tilde{S}$ for the sequential game, given by positive functionals $\phi_{xa}$ and POVM elements $\{\tilde{B}_{yb}\}$, such that
$\lim_{k\to\infty}\operatorname{tr}\left(\sigma_{xa}^{\lambda_{\ell_{k}}}\,P(\{B^{\lambda_{\ell_{k}}}_{yb}\})\right)=\phi_{xa}(P(\{\tilde{B}_{yb}\}))$ (6.5)
and
$\omega(\mathcal{G}_{\textnormal{seq}},\tilde{S})=\lim_{k\to\infty}\omega_{\lambda_{\ell_{k}}}(\mathcal{G}_{\textnormal{comp}},S)=\lim_{\ell\to\infty}\omega_{\lambda_{\ell}}(\mathcal{G}_{\textnormal{comp}},S)=\omega_{qc}(\mathcal{G}).$
Now we can apply Lemma 6.4 to deduce Eq. 6.4 from Eq. 6.5. ∎
@@ Acknowledgements
We thank William Slofstra for bringing [34] to our attention and Victoria Wright and Lewis Wooltorton for drawing our attention to [46]. We also thank Victoria Wright for bringing [8, 43] to our attention and for pointing out that our characterization of commuting operator correlations resolves an open problem in [46]. Finally, we thank Michal Banacki for bringing his independent work [1] to our attention after a preprint of our paper had appeared online. CP thanks the hospitality at RUB for a productive visit. MW acknowledges the Simons Institute for the Theory of Computing at UC Berkeley for its hospitality and support. CP acknowledges funding support from the Natural Sciences and Engineering Research Council of Canada (NSERC). AK, GM, SS and MW acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2092 CASA - 390781972. GM is also supported by the European Research Council through an ERC Starting Grant (Grant agreement No. 101077455, ObfusQation). MW is supported by the European Research Council through an ERC Starting Grant (Grant agreement No. 101040907, SYMOPTIC), by the NWO through grant OCENW.KLEIN.267, and by the BMBF through projects Quantum Methods and Benchmarks for Resource Allocation (QuBRA) and Quantum Algorithm Solver Toolkit (QuSol).